OpenStudy (raiders88):

evaluate the function for x=2: f(x) = 7x-5

OpenStudy (texaschic101):

f(x) = 7x - 5 f(2) = 7(2) - 5 f(2) = 14 - 5 f(2) = 9 understand ?

OpenStudy (raiders88):

uh huh

OpenStudy (texaschic101):

basically, your just subbing in 2 for x

OpenStudy (raiders88):

i have two more if u will help

OpenStudy (texaschic101):

I can try...but I am not the best at these...just the simple ones

OpenStudy (raiders88):

candy is on sale for $0.75 each. you have a coupon for $0.25 off your total purchase. write a function rule for the cost of n pieces of candy

OpenStudy (texaschic101):

f(n) = 0.75n - 0.25

OpenStudy (raiders88):

k last one

OpenStudy (raiders88):

the ordered pairs (0,0),(2,4), (4,8), and (6,12) represent a function. what is the rule for this function

OpenStudy (texaschic101):

Not too sure.....all I know is the x terms add 2 and the y terms add 4...but I do not know how to write this as a function...sorry

OpenStudy (raiders88):

its ok

OpenStudy (raiders88):

@IrishBoy123